# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ 5 Hierarchical Task Networks
Suppose there is a multicolored clique of size $k$. Decompose according to the above interpretation. We obtain a primitive network with a total order on the tasks. The corresponding task sequence is executable in $tn$ from $s_{0}$, and as it executes all tasks including $t_{g}$, it yields a yes-instance for Plan Existence, Action Executability, and State Reachability.
@@
For the constructed instances, ${C_{d}}={C_{s}}=1$ and ${C_{c}}=2$. For the two other combinations of parameters for which the statement holds, we employ the same approaches we used to adapt Theorem 17. In particular, to prove the statement for ${C_{c}}=2$, ${C_{d}}=2$, and ${C_{\#}}=1$, we let the initial network contain a single compound task that is decomposed into a network with the above structure. To prove the statement for ${C_{c}}=2$, ${C_{\#}}=1$, and ${C_{s}}=2$, we let the initial network only contain the compound task $V_{1}$. Then, we adapt all the decomposition methods. For each method $(c,tn_{m})\in M$, the network $tn_{m}$ consisted of a single primitive task $t_{m}$. Let $t^{\prime}$ denote the task identifier that followed $c$ in the total order in the initial network. Instead of decomposing into a single primitive task $t_{m}$, let the method now decompose into a network with tasks $t_{m}$ and $t^{\prime}$ with $(t_{m},t^{\prime})\in\prec^{+}$. Note that, in this construction, exactly the same set of primitive networks can be created as in the original construction, but now ${C_{\#}}=1$ and ${C_{d}}$ is not bounded. ∎
6 A Parameterized Analysis of HTNs
In this final section, we ask whether—and under which parameterizations—the fundamental problems on HTNs admit fixed-parameter algorithms.
We begin by noting that while there is by now a very broad range of graph-theoretic parameters which can typically be used to achieve fixed-parameter tractability, none of the “usual suspects” in that regard will help when dealing with any of the four problems of interest here. In particular, Theorems 2 and 5 immediately exclude the algorithmic application of not only treewidth Robertson and Seymour (1986)—the by far most commonly used structural graph parameter—but also treedepth Nesetril and de Mendez (2012), the feedback edge number Ganian and Korchemna (2021) and even the broad family of analogous measures defined specifically on directed graphs Ganian et al. (2016).
Moreover, while the positive results obtained in Section 4 establish that our problems of interest are polynomial-time solvable when the generalized partial order width is fixed, none of the problems is fixed-parameter tractable when parameterized by the generalized partial order width (or even by the “usual” partial order width). Indeed, since the Shuffle Product problem used in the reductions underlying the proofs of Theorems 3 and 4 is also known to be W[2]-hard when parameterized by the width $w$ (van Bevern et al., 2016, Theorem 2.4), it turns out that the proofs of these two theorems also immediately establish the following:
Corollary 21.
Plan Validation on primitive networks where $\prec^{+}$ partitions $T$ into $w$ chains such that there are no dependencies between the chains is W[2]-hard parameterized by $w$, even if $|A|=2$ and $\operatorname{prec}(a)=\operatorname{del}(a)=\operatorname{add}(a)=\emptyset$ for all actions $a\in A$.
Corollary 22.
State Reachability and Plan Existence on primitive networks where $\prec^{+}$ partitions $T$ into $w$ chains are W[2]-hard parameterized by $w$, even if $|A|=5$ and the state transition graph has $k=4$ states.
Fixed-Parameter Tractability via Vertex Cover.
In this subsection, we contrast the above lower bounds with a fixed-parameter algorithm that exploits a different parameterization of the network: we show that Plan Validation is fixed-parameter tractable when parameterized by the vertex cover number of $G_{\prec}$. Recall that a vertex cover of $G_{\prec}$ is a subset of tasks $V\subseteq T$ such that, for all $t\prec t^{\prime}$, $t\in V$ and/or $t^{\prime}\in V$; the vertex cover number is the minimum size of a vertex cover of $G_{\prec}$. We remark that, while the vertex cover number has also been used to obtain fixed-parameter algorithms for many other problems Bhore et al. (2020); Balko et al. (2022); Blazej et al. (2023), the techniques from previous works do not easily translate to HTN planning problems. Instead, our algorithms make use of a delicate branching routine whose correctness requires a careful analysis of the state space.
Theorem 23.
Plan Validation is fixed-parameter tractable parameterized by the vertex cover number of $G_{\prec}$.
Proof.
@@ 6 A Parameterized Analysis of HTNs
By definition, $p$ is executable if and only if the actions in the plan have their preconditions fulfilled and each $t\in T$ can be assigned a unique index in the sequence such that the assignment according to $\alpha$ matches the plan and the imposed order does not conflict with $\prec^{+}$. Observe that simulating the action sequence suffices to check whether all preconditions are met at the respective step, so it suffices to handle the assignment of tasks to indices. We branch over all possible orderings of the vertices in $V$. Orderings that do not respect an edge in $\prec^{+}\cap(V\times V)$ are discarded immediately. For $i\in[|V|]$, let $v_{i}$ denote the $i^{\text{th}}$ task in the ordering. Then, for all $t\in T\setminus V$, let the priority of $t$ be defined as $\psi(t)=\min_{(t,v_{i})\in\prec}i$ and $\psi(t)=\infty$ if there is no such edge. For all $v_{i}\in V$, let $\psi(v_{i})=i+0.5$. We now iterate through $p$ and assign tasks to each index in a greedy manner. Assume we have already assigned tasks to the first $i-1$ actions for some $i\in[|T|]$. Then, we say that $t\in T$ is executable as the $i^{\text{th}}$ action if $\alpha(t)=p[i]$, it is not yet assigned an index, and all its predecessors in $\prec^{+}$ are assigned to indices less than $i$. Among all executable tasks at the index $i$, we choose the one with the lowest priority according to $\psi$. If there is no executable task for some index, then we return that this branch does not yield a yes-instance. The algorithm returns that it is a yes-instance if any branch manages to assign all the tasks to $p$, and otherwise, returns that it is a no-instance.
@@
The algorithm branches over the $|V|!$ possible orderings of tasks in the vertex cover. For each ordering, the priorities can be computed in $\mathcal{O}\!\left(|\prec|+|T|\right)$ time. By employing a Hollow Heap Hansen et al. (2017) to implement a priority queue, iterating through the sequence, extracting the minimum priority executable task, and inserting all then executable tasks to the queue takes $\mathcal{O}\!\left(|\prec|+|T|\log|T|\right)$ total time, yielding a runtime of $\mathcal{O}\!\left(|V|!(|\prec|+|T|\log|T|)\right)$ for the entire algorithm. ∎
By Theorem 1, tackling the remaining problems requires additional restrictions on the state space. We show that here, too, we can achieve tractability, albeit the arguments are even more involved than in the previous case.
Theorem 24.
Plan Validation and Action Executability are fixed-parameter tractable when parameterized by the vertex cover number of $G_{\prec}$ plus the number of states in the state transition graph.
Proof.
Let the input instance consist of the primitive network $tn=(T,\prec^{+},\alpha)$, an initial state $s_{0}$, and either a target state $s_{g}$ (for State Reachability) or an action set $S$ (for Action Executability). For Action Executability, assume that Reduction Rule R0 cannot be applied (as otherwise we can solve the instance directly). Begin by computing a minimum-size vertex cover $V$ of $G_{\prec}$ by the classical fixed-parameter algorithm Downey and Fellows (2013), and let $K$ denote the number of states in the state transition graph ${\cal G}$.
@@
For each branch, we define a strong equivalence relation on the tasks. For all $t\in T\setminus V$, the admissible interval is defined as $[\max{\{i\mid v_{i}\prec t\}},\min{\{i\mid t\prec v_{i}\}}-1]$. [^2] If the first set is empty, then the interval starts at 0. If the second set is empty, then it ends at $|V^{\prime}|$. Then, $t_{1},t_{2}\in T$ are strongly equivalent if and only if $\alpha(t_{1})$ and $\alpha(t_{2})$ are equivalent and have the same admissible interval. There are at most $E=(k+1)^{k}\cdot(|V^{\prime}|^{2}+1)$ strong equivalence classes. Define the augmented state transition graph ${\cal G}^{\prime}$ over the same set of vertices as the regular state transition graph, but let it have no edges for actions. Instead, let there be an edge $(e,s,s^{\prime})$ for all strong equivalence classes $e$ and states $s,s^{\prime}$ such that $e$ transforms $s$ into $s^{\prime}$. For each such edge, let $\beta((e,s,s^{\prime}))=(s,s^{\prime})$.
@@
Further, let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are considered distinct if and only if they do not contain the exact same vertices, i.e., cycles obtained from other cycles by inverting their directions are not counted as different cycles. Then, $|C|\leq(2E)^{k}$. For each $0\leq i\leq|V^{\prime}|$, we branch over all possible sets $C_{i}\subseteq C$, the cycles that are walked as a detour from path $W_{i}$. This splits each prior branch into ${\mathcal{O}\!\left((2^{|C|})^{|V^{\prime}|+1}\right)}\leq{\mathcal{O}\!\left(2^{(2E)^{k}\cdot(|V^{\prime}|+1)}\right)}$ new branches. Thus, the total number of branches is
@@
We build a solution plan that walks each of the paths $W_{i}$, $0\leq i\leq|V^{\prime}|$, and additionally walks through the associated cycles. The exact number of times a cycle is walked is determined by an ILP. This ILP has a variable $x_{c}$ for every $c\in C$. For each strong equivalence class $e$, let $\max{{}^{\prime}}(e)$ denote the number of tasks in $T\setminus V$ in equivalence class $e$. For all $c\in C$, let $c(e)$ denote the number of occurrences of $e$ in $c$. Further, let $w_{i}(e)$ denote the number of occurrences of $e$ in $W_{i}$ for $0\leq i\leq|V^{\prime}|$. Note that, in any plan that is a solution to the problem instance, two strongly equivalent tasks can be swapped and the plan will still be valid. Thus, to ensure no task in $T\setminus V$ is executed more than once, for each equivalence class $e$ we have a constraint
@@
Only for Action Executability instances, for each action equivalence class $e_{a}$, let $\min^{\prime}(e_{a})$ be the number of occurrences of $e_{a}$ in $S$, and let $w_{i}^{\prime}(e_{a})$ and $c_{i}^{\prime}(e_{a})$ be the number occurrences of $e_{a}$ in path $W_{i}$ and cycle $c_{i}$, respectively. Then, add a constraint
@@
If the ILP has a solution, then construct a sequence of equivalence classes as follows. For each $0\leq i\leq|V^{\prime}|$, take any cycle $c\in C_{i}$ that shares a vertex $s$ with the current path $W_{i}$. Delete $c$ from $C_{i}$ and augment $W_{i}$ by inserting cycle $c$ at that position once. If there is no more $0\leq j\leq|V|^{\prime}$, $j\neq i$, such that $c\in C_{j}$, instead add as many repetitions of $c$ to $W_{i}$ as the number of times that cycle $c$ was added to the path exactly $x_{c}$ times in total. As the cycles end in the same state they started in, the resulting action sequence is executable. Repeat until $C_{i}$ is empty. As $C_{i}$ and $W_{i}$ form a connected component, all cycles in $C_{i}$ can be added this way. Construct a solution task sequence by concatenating the augmented paths $W_{0},\ldots,W_{|V^{\prime}|}$ and, for a step by a certain equivalence class, employ any task of that class. For Action Executability, choose tasks with a not yet covered action in $S$ whenever possible. Right before each path $W_{i}$,$i\in[|V^{\prime}|]$, execute task $v_{i}$. Observe that this way the preconditions of each task are fulfilled when it is about to be executed. Further, branches in which there are tasks that are executed before their predecessors or after their successors in $\prec^{+}$ are already discarded. The ILP ensures that each equivalence class $e$ appears at most $\max{{}^{\prime}}(e)$ times. As there are $\max{{}^{\prime}}(e)$ tasks of that class and they can be used interchangeably, the sequence is a valid construction. For State Reachability, only paths that end in $s_{g}$ are considered. For Action Executability, the additional set of constraints ensures that $S$ is covered as there are enough tasks of each action and tasks from the same strong equivalence class can be used interchangeably. Thus, the given instance is a yes-instance.
@@
For the other direction, suppose there is a task sequence solving the instance. There is a branch of the algorithm that considers exactly the set $V^{\prime}$ of vertex cover tasks used in the sequence in the order $v_{1},\ldots,v_{|V|}$ and executes each task $v_{i}$ from state $s_{i}$. Split the sequence at the vertex cover tasks into subwalks $W_{0}^{*},\ldots,W_{|V|}^{*}$. Then, for all $0\leq i\leq|V^{\prime}|$, consider an initially empty set of cycles $C^{*}_{i}$. Find any substring $W_{i}^{*}$ that describes a simple edge cycle in ${\cal G}^{\prime}$. Add this cycle to $C^{*}_{i}$ (disregarding direction and starting vertex) and remove the respective substring from $W_{i}^{*}$. Repeat until no more such cycles exist. Observe that each $W_{i}^{*}$ then describes a simple edge path in ${\cal G}^{\prime}$, and all $C^{*}_{i}$ exclusively contain simple edge cycles in ${\cal G}^{\prime}$. We write $W_{i}^{\prime}$ for these reduced $W_{i}^{*}$. There is a branch of the algorithm such that $W_{i}=W_{i}^{\prime}$ and $C_{i}=C_{i}^{*}$ for all $0\leq i\leq|V^{\prime}|$. This branch is not discarded because the task sequence is valid, so $\operatorname{prec}(v_{i})\subseteq s_{i}$ for all $i\in[|V^{\prime}|]$, and the paths $W_{i}$ and cycles in $C_{i}$ are such that they do not depend on any vertex cover task after $v_{i}$, and no vertex cover task up to $v_{i}$ depends on them. Further, as the cycles in $C_{i}^{*}$ were iteratively removed from the path $W_{i}^{*}$, the paths and cycles in $W_{i}$ and $C_{i}$ form a connected component in ${\cal G}^{\prime}$. For each cycle $c\in C$, let $x_{c}^{*}$ be the total number of times $c$ has been removed from paths $W_{i}^{*}$. Then, the ILP has a solution if $x_{c}=x_{c}^{*}$ for all $c\in C$ as follows. As the $x_{c}^{*}$ are computed from a valid sequence, summing over all occurrences of an equivalence class $e$ in paths $W_{i}$ and cycles in $C_{i}$ does not exceed $\max{{}^{\prime}}(e)$. Further, by construction, $x_{c}^{*}$ is at least the number of sets $C_{i}$ with $c\in C_{i}$. In the case of a valid Action Executability solution, the number of occurrences of each action equivalence class $e_{a}$ is at least $\min{{}^{\prime}}(e_{a})$, and thus, the additional constraints are fulfilled. Hence, the ILP has a solution and the algorithm correctly decides for a yes-instance.
@@
Employing reduction rule R0 and building the augmented state transition graph takes $\mathcal{O}\!\left(|A|k+|T|\right)$ time. To decide whether a branch is discarded takes $\mathcal{O}\!\left(|\prec|+|T|+|V||C|^{2}k\right)$ time. Computing $c(e)$, $w_{i}(e)$, $\max^{\prime}(e)$, $c^{\prime}(e_{a})$, $w_{i}^{\prime}(e_{a})$, $\min^{\prime}(e_{a})$, and |{$|{\{0\leq i\leq|V^{\prime}|\mid c\in C_{i}\}}|$ takes $\mathcal{O}\!\left(|T|+(|C|+|V|+1)k+|C|\cdot(|V|+1)\right)$ time. The ILP has $|C|$ variables and at most $E+|C|+(k+1)^{k}={\mathcal{O}\!\left(E+|C|\right)}$ constraints, in which the absolute values of all coefficients are bounded by $\max{\{k,|T|\}}$. As noted before, an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. Thus, the ILP is solved in

### v2

@@ Appendix A Appendix: Proof Details
Suppose there is a multicolored clique of size $k$. Decompose according to the above interpretation. We get a primitive network with a total order on the tasks. The corresponding action sequence is executable from $s_{0}$ and executes all tasks including $t_{g}$, witnessing a yes-instance for Plan Existence, Action Executability, and State Reachability.
@@
Note that for the constructed instances, ${C_{d}}={C_{s}}=1$ and ${C_{c}}=2$, proving the case of non-constant ${C_{\#}}$. For non-constant ${C_{s}}$, we let the initial network contain a single compound task that is decomposed into a network with the above structure. Then, ${C_{c}}=2$, ${C_{d}}=2$, and ${C_{\#}}=1$. For non-constant ${C_{d}}$, we let the initial network only contain the compound task $V_{1}$. Then, we adapt all the decomposition methods. For each method $(t,tn_{m})\in M$, the network $tn_{m}$ consists of a single primitive task $t_{m}$. Let $t^{\prime}$ denote the task identifier that followed $t$ in the total order in the initial network. Instead of decomposing into a single primitive task $t_{m}$, let the method decompose into a network with tasks $t_{m}$ and $t^{\prime}$ with $(t_{m},t^{\prime})\in\prec^{+}$. Note that the exact same set of primitive networks can be created as in the original construction, but now ${C_{d}}$ is non-constant, ${C_{c}}={C_{s}}=2$, and ${C_{\#}}=1$.
All initial task networks have (generalized) partial order width 1 or 0 as they consist of a chain or a single vertex. ∎
Proof of Theorem 10.
@@
We have a yes-instance if the input plan $p$ is executable in $s_{0}$ and each $t\in T$ can be assigned a unique index in the sequence such that the assignment according to $\alpha$ matches the plan and the imposed order does not conflict with $\prec^{+}$. Observe that simulating the action sequence suffices to check whether all preconditions are met at the respective step, so it remains to handle the assignment of tasks to indices. We branch over all possible orderings of the vertices in $V$. Orderings that do not respect an edge in $\prec^{+}\cap(V\times V)$ are discarded immediately. For $i\in[|V|]$, let $v_{i}$ denote the $i^{\text{th}}$ task in the ordering. Then, for all $t\in T\setminus V$, let the priority of $t$ be defined as $\psi(t)=\min_{(t,v_{i})\in\prec}i$ and $\psi(t)=\infty$ if there is no such edge. For all $v_{i}\in V$, let $\psi(v_{i})=i+0.5$. We now iterate through $p$ and assign tasks to each index in a greedy manner. Assume we have already assigned tasks to the first $i-1$ actions for some $i\in[|T|]$. Then, we say that $t\in T$ is executable as the $i^{\text{th}}$ action if $\alpha(t)=p[i]$, it is not yet assigned an index, and all its predecessors in $\prec^{+}$ are assigned to indices less than $i$. Among all executable tasks at the index $i$, we choose the one with the lowest priority according to $\psi$. If there is no executable task for some index, then we return that this branch does not yield a yes-instance. The algorithm returns that it is a yes-instance if any branch manages to assign all tasks to $p$, and rejects otherwise.
@@
The algorithm branches over the $|V|!$ possible orderings of tasks in the vertex cover. For each ordering, the priorities can be computed in $\mathcal{O}\!\left(|{\prec}|+|T|\right)$ time. Using a Hollow Heap Hansen et al. [2017] to implement a priority queue, iterating through the sequence, extracting the minimum priority executable task, and inserting all newly executable tasks to the queue takes $\mathcal{O}\!\left(|{\prec}|+|T|\log|T|\right)$ time, yielding a runtime of $\mathcal{O}\!\left(|V|!(|{\prec}|+|T|\log|T|)\right)$ for the entire algorithm. ∎
Proof of Theorem 11.
Let the input instance consist of the primitive network $tn=(T,\prec^{+},\alpha)$, an initial state $s_{0}$, and either a target state $s_{g}$ (for State Reachability) or an action set $S$ (for Action Executability). For Action Executability, assume that Reduction Rule R0 cannot be applied (as otherwise we can solve the instance directly). First use a fixed-parameter algorithm to compute a minimum-size vertex cover $V$ of $G_{\prec}$ Downey and Fellows [2013], and let $k$ denote the number of states in the state transition graph ${\cal G}$.
@@
For each branch, we define a strong equivalence relation on the tasks. For all $t\in T\setminus V$, the admissible interval is defined as $[\max{\{i\mid v_{i}\prec t\}},\min{\{i\mid t\prec v_{i}\}}-1]$. [^2] If the first set is empty, then the interval starts at 0. If the second set is empty, then it ends at $|V^{\prime}|$. Then, $t_{1},t_{2}\in T$ are strongly equivalent if and only if $\alpha(t_{1})$ and $\alpha(t_{2})$ are equivalent and have the same admissible interval. There are at most $E=(k+1)^{k}\cdot(|V^{\prime}|^{2}+1)$ strong equivalence classes. Define the augmented state transition graph ${\cal G}^{\prime}$ over the same set of vertices as the regular state transition graph, but let it have no edges for actions. Instead, let there be an edge with label $e$ from $s$ to $s^{\prime}$ for all strong equivalence classes $e$ and states $s,s^{\prime}$ such that $e$ transforms $s$ into $s^{\prime}$.
@@
Further, let $C$ be the set of distinct simple edge cycles in ${\cal G}$, where two cycles are distinct if and only if they do not contain the same vertices. Then, $|C|\leq(2E)^{k}$. For each $0\leq i\leq|V^{\prime}|$, we branch over all possible sets $C_{i}\subseteq C$, the cycles that are walked as a detour from path $W_{i}$. This splits each prior branch into ${\mathcal{O}\!\left((2^{|C|})^{|V^{\prime}|+1}\right)}\leq{\mathcal{O}\!\left(2^{(2E)^{k}\cdot(|V^{\prime}|+1)}\right)}$ new branches. Thus, the total number of branches is
@@
We build a solution plan that walks each of the paths $W_{i}$, $0\leq i\leq|V^{\prime}|$, and additionally walks through the associated cycles. The number of times a cycle is walked is determined by an ILP with a variable $x_{c}$ for every $c\in C$. For each strong equivalence class $e$, let ${\max}^{\prime}(e)$ denote the number of tasks in $T\setminus V$ in equivalence class $e$. For all $c\in C$, let $c(e)$ denote the number of occurrences of $e$ in $c$. Further, let $w_{i}(e)$ denote the number of occurrences of $e$ in $W_{i}$ for $0\leq i\leq|V^{\prime}|$. Note that, in any plan that is a solution to the problem instance, two strongly equivalent tasks can be swapped and the plan will still be valid. Thus, to ensure no task in $T\setminus V$ is executed more than once, for each equivalence class $e$ add a constraint
@@
For Action Executability, for each action equivalence class $e_{a}$, let $\min^{\prime}(e_{a})$ be the number of occurrences of $e_{a}$ in $S$, and let $w_{i}^{\prime}(e_{a})$ and $c_{i}^{\prime}(e_{a})$ be the number occurrences of $e_{a}$ in path $W_{i}$ and cycle $c_{i}$, respectively. Then, add a constraint
@@
If the ILP has a solution, then construct a sequence of equivalence classes as follows. For each $0\leq i\leq|V^{\prime}|$, take any cycle $c\in C_{i}$ that shares a vertex $s$ with the current path $W_{i}$. Delete $c$ from $C_{i}$ and augment $W_{i}$ by inserting cycle $c$ at that position once. If there is no more $0\leq j\leq|V|^{\prime}$, $j\neq i$, such that $c\in C_{j}$, instead add as many repetitions of $c$ to $W_{i}$ as the number of times that cycle $c$ was added to the path precisely $x_{c}$ times in total. As the cycles end in the same state they started in, the resulting action sequence is executable. Repeat until $C_{i}$ is empty. As $C_{i}$ and $W_{i}$ form a connected component, all cycles in $C_{i}$ can be added this way. Construct a solution task sequence by concatenating the augmented paths $W_{0},\ldots,W_{|V^{\prime}|}$ and, for a step by a certain equivalence class, employ any task of that class. For Action Executability, choose tasks with a not yet covered action in $S$ whenever possible. Right before each path $W_{i}$,$i\in[|V^{\prime}|]$, execute task $v_{i}$. Observe that this way the preconditions of each task are fulfilled when it is about to be executed. Further, branches in which there are tasks that are executed before their predecessors or after their successors in $\prec^{+}$ are already discarded. The ILP ensures that each equivalence class $e$ appears at most $\max{{}^{\prime}}(e)$ times. As there are $\max{{}^{\prime}}(e)$ tasks of that class and they can be used interchangeably, the sequence is a valid construction. For State Reachability, only paths that end in $s_{g}$ are considered. For Action Executability, the additional set of constraints ensures that $S$ is covered as there are enough tasks of each action and tasks from the same strong equivalence class can be used interchangeably. Thus, the given instance is a yes-instance.
@@
For the other direction, suppose there is a task sequence witnessing a yes-instance. There is a branch of the algorithm that considers exactly the set $V^{\prime}$ of vertex cover tasks used in the sequence in the order $v_{1},\ldots,v_{|V|}$ and executes each task $v_{i}$ from state $s_{i}$. Split the sequence at the vertex cover tasks into subwalks $W_{0}^{*},\ldots,W_{|V|}^{*}$. Then, for all $0\leq i\leq|V^{\prime}|$, consider an initially empty set of cycles $C^{*}_{i}$. Find any substring $W_{i}^{*}$ that describes a simple edge cycle in ${\cal G}^{\prime}$. Add this cycle to $C^{*}_{i}$ (disregarding direction and starting vertex) and remove the respective substring from $W_{i}^{*}$. Repeat until no more such cycles exist. Observe that each $W_{i}^{*}$ then describes a simple edge path in ${\cal G}^{\prime}$, and all $C^{*}_{i}$ exclusively contain simple edge cycles in ${\cal G}^{\prime}$. We write $W_{i}^{\prime}$ for these reduced $W_{i}^{*}$. There is a branch of the algorithm such that $W_{i}=W_{i}^{\prime}$ and $C_{i}=C_{i}^{*}$ for all $0\leq i\leq|V^{\prime}|$. This branch is not discarded because the task sequence is valid, so $\operatorname{prec}(v_{i})\subseteq s_{i}$ for all $i\in[|V^{\prime}|]$, and the paths $W_{i}$ and cycles in $C_{i}$ are such that they do not depend on any vertex cover task after $v_{i}$, and no vertex cover task up to $v_{i}$ depends on them. Further, as the cycles in $C_{i}^{*}$ were iteratively removed from the path $W_{i}^{*}$, the paths and cycles in $W_{i}$ and $C_{i}$ form a connected component in ${\cal G}^{\prime}$. For each cycle $c\in C$, let $x_{c}^{*}$ be the total number of times $c$ has been removed from paths $W_{i}^{*}$. Then, the ILP has a solution if $x_{c}=x_{c}^{*}$ for all $c\in C$ as follows. As the $x_{c}^{*}$ are computed from a valid sequence, summing over all occurrences of an equivalence class $e$ in paths $W_{i}$ and cycles in $C_{i}$ does not exceed $\max{{}^{\prime}}(e)$. Further, by construction, $x_{c}^{*}$ is at least the number of sets $C_{i}$ with $c\in C_{i}$. In the case of an Action Executability solution, the number of occurrences of each action equivalence class $e_{a}$ is at least ${\min}^{\prime}(e_{a})$, satisfying the additional constraints. Hence, the ILP has a solution and the algorithm correctly decides for a yes-instance.
@@
Applying reduction rule R0 and building the augmented state transition graph takes $\mathcal{O}\!\left(|A|k+|T|\right)$ time. Deciding whether to discard a branch takes $\mathcal{O}\!\left(|{\prec}|+|T|+|V||C|^{2}k\right)$ time. Computing $c(e)$, $w_{i}(e)$, $\max^{\prime}(e)$, $c^{\prime}(e_{a})$, $w_{i}^{\prime}(e_{a})$, $\min^{\prime}(e_{a})$, and |{$|{\{0\leq i\leq|V^{\prime}|\mid c\in C_{i}\}}|$ takes $\mathcal{O}(|T|+(|C|+|V|+1)k+|C|\cdot(|V|+1))$ time. The ILP has $|C|$ variables and at most $E+|C|+(k+1)^{k}={\mathcal{O}\!\left(E+|C|\right)}$ constraints, in which the absolute values of all coefficients are bounded by $\max{\{k,|T|\}}$. As noted before, an ILP instance ${\cal I}_{\text{ILP}}$ with $n$ variables can be solved in $\mathcal{O}\!\left(n^{O(n)}\cdot|{\cal I}_{\text{ILP}}|\right)$ time. Thus, the ILP is solved in